Could I please have some help with this word problem? I started it, but am now lost. Thanks so much.
A garden area is 30 ft. long and 20 ft. wide. A path of uniform width is set around the edge. If the remaining garden area is 400ft.^2, what is the width of the path?
I began by using the equation (20+2x)(30+2x)=400, but I was not able to get a real number.


Answer by checkley75(3666)

(20+2X)(30+2X)=400
600+60X+40X+4X^2=400
4X^2+100X+600-400=0
4X^2+100X+200=0
X^2+25X+50=0 USING THE QUADRATIXC EQUATION WE GET
X=(-25+-SQRT[25^2-4*1*50])/2*1
X=(-25+-SQRT[625-200])/2
X=(-25+-SQRT425)/2
X=(-25+-20.62)/2
X=(-25+20.62)/2
X=-4.38/2
X=-2.19 ANSWER.
(20+2*-2.19)(30+2*-2.19)=400
(20-4.38)(30-4.38)=400
15.62*25.62=400
400=400